OpenStudy (anonymous):

Please Help! Simplify the radical expression. (-)square root of 7y^6/16x^4

OpenStudy (ranga):

you ca take the square root of individual terms in the numerator and the denominator. sqrt(7) cannot be simplified any further and so it will remain as sqrt(7) What is sqrt(y^6) ? (remember: sqrt(y^6) = (y^6)^(1/2) = ?

OpenStudy (anonymous):

y^3?

ganeshie8 (ganeshie8):

yes, wat about denominator

ganeshie8 (ganeshie8):

\(\huge - \sqrt{\frac{7y^6}{16x^4}}\)

ganeshie8 (ganeshie8):

\(\huge - \frac{\sqrt{7y^6}}{\sqrt{16x^4}}\)

ganeshie8 (ganeshie8):

\(\huge - \frac{y^3\sqrt{7}}{\sqrt{16x^4}}\)

ganeshie8 (ganeshie8):

simplify denominator also

OpenStudy (isaiah.feynman):

For denominator use this rule